# DynaSearcher: Dynamic Knowledge Graph Augmented Search Agent via Multi-Reward Reinforcement Learning

### v1

@@ Abstract
Multi-step agentic retrieval systems based on large language models (LLMs) have demonstrated remarkable performance in complex information search tasks. However, in practical applications, these systems are limited by factually inconsistent intermediate queries and inefficient search trajectories, which can cause reasoning deviations and redundant computations. To address these issues, we propose DynaSearcher, an innovative search agent enhanced by dynamic knowledge graphs and multi-reward reinforcement learning (RL). Specifically, our system leverages knowledge graphs as external structured knowledge to guide the search process by explicitly modeling entity relationships, thereby ensuring factual consistency in intermediate queries and mitigating biases from irrelevant information. Furthermore, we employ a multi-reward RL framework for fine-grained control over training objectives such as retrieval accuracy, efficiency, and response quality. This framework promotes the generation of high-quality intermediate queries and comprehensive final answers, while discouraging unnecessary exploration and minimizing information omissions or redundancy. Experimental results demonstrate that our approach achieves state-of-the-art answer accuracy on six multi-hop question answering datasets and exhibits strong generalization and robustness across diverse retrieval environments and larger-scale models, highlighting its broad applicability.
@@
Chuzhan Hao Wenfeng Feng Yuewei Zhang Hao Wang††thanks: Corresponding author Alibaba Cloud Computing {haochuzhan.hcz,wenfeng.fwf,liyou.zyw}@alibaba-inc.com cashenry@126.com
@@ 3.1 Rollout with Search Tools
Doc Search Tool. We employ two types of document search tools: a locally deployed vector-based retrieval service and the web search provided by Tavily111https://tavily.com. Both tools return text-based retrieval results. In the experiments, we fix a top-k value (e.g., 5) to control the number of retrieved documents.
@@ 3.2.2 Gain and Penalty Reward
where denotes the trainable policy model, is a fixed reference model, represents the overall reward function, and denotes the KL divergence. Here, are sampled from the dataset , and denote the output sequence interleaving reasoning steps with search engine retrievals. Since the retrieved documents are not generated by the policy model, we mask the retrieval results during loss calculation to prevent the training policy from being biased.
@@ 4.1 Experimental Settings
Baselines and Training Details. In our experiments, in addition to comparing with state-of-the-art LLMs such as DeepSeek-R1-0528, Qwen3-235B-A22B, GPT-4.1-0414, o4-mini-0416, and Gemini-2.5-Pro-0325 (as shown in Table 2), we also benchmark against advanced RAG methods Shao et al. (2023); Trivedi et al. (2022a); Li et al. (2025a) and RL-based agentic search models Jin et al. (2025); Chen et al. (2025); Song et al. (2025); Zheng et al. (2025); Wang et al. (2025). These experiments are primarily based on the Qwen2.5 models Qwen et al. (2025), where Qwen2.5-7B and Qwen2.5-32B refer to their respective Instruct models. All training-based models are derived from their corresponding open-source models. The training data of DynaSearcher consist of the stage-2 data from Song et al. (2025) and 8,000 randomly sampled instances from Musique. The RL training epoch is set to 1, with a train batch size of 16 and a learning rate of 1e-6. The KL divergence coefficient set to 1e-3. Each data sample undergoes 8 rollouts during training. We utilize FSDP Zhao et al. (2023) and vLLM222https://github.com/vllm-project/vllm in VeRL333https://github.com/volcengine/verl framework, with a sampling temperature of 1.0, top-p of 0.95 and a maximum response length of 8192. For our information gain and penalty reward, we set to 0.5, to 0.9, and to -0.2.

### v2

@@ Abstract
Multi-step agentic retrieval systems based on large language models (LLMs) have demonstrated remarkable performance in complex information search tasks. However, in practical applications, these systems are limited by factually inconsistent intermediate queries and inefficient search trajectories, which can cause reasoning deviations and redundant computations. To address these issues, we propose DynaSearcher, an innovative search agent enhanced by dynamic knowledge graphs and multi-reward reinforcement learning (RL). Specifically, our system leverages knowledge graphs as external structured knowledge to guide the search process by explicitly modeling entity relationships, thereby ensuring factual consistency in intermediate queries and mitigating biases from irrelevant information. Furthermore, we employ a multi-reward RL framework for fine-grained control over training objectives such as retrieval accuracy, efficiency, and response quality. This framework promotes the generation of high-quality intermediate queries and comprehensive final answers, while discouraging unnecessary exploration and minimizing information omissions or redundancy. Experimental results demonstrate that our approach achieves state-of-the-art answer accuracy on six multi-hop question answering datasets and exhibits strong generalization and robustness across diverse retrieval environments and larger-scale models, highlighting its broad applicability111The model is available at https://modelscope.cn/collections/DynaSearcher-a00139d1ef2542..
@@
Chuzhan Hao Wenfeng Feng Yuewei Zhang Hao Wang††thanks: Corresponding author. Alibaba Cloud Computing haochuzhan.hcz@alibaba-inc.com, cashenry@126.com
@@ 3.1 Rollout with Search Tools
Doc Search Tool. We employ two types of document search tools: a locally deployed vector-based retrieval service and the web search provided by Tavily222https://tavily.com.. Both tools return text-based retrieval results. In the experiments, we fix a top-k value (e.g., 5) to control the number of retrieved documents.
@@ 3.2.2 Gain and Penalty Reward
where denotes the trainable policy model, is a fixed reference model, represents the overall advantage function, and denotes the KL divergence. Here, are sampled from the dataset , and denote the output sequence interleaving reasoning steps with search engine retrievals. Since the retrieved documents are not generated by the policy model, we mask the retrieval results during loss calculation to prevent the training policy from being biased.
@@ 4.1 Experimental Settings
Baselines and Training Details. In our experiments, in addition to comparing with state-of-the-art LLMs such as DeepSeek-R1-0528, Qwen3-235B-A22B, GPT-4.1-0414, o4-mini-0416, and Gemini-2.5-Pro-0325 (as shown in Table 2), we also benchmark against advanced RAG methods Shao et al. (2023); Trivedi et al. (2022a); Li et al. (2025a) and RL-based agentic search models Jin et al. (2025); Chen et al. (2025); Song et al. (2025); Zheng et al. (2025); Wang et al. (2025). These experiments are primarily based on the Qwen2.5 models Qwen et al. (2025), where Qwen2.5-7B and Qwen2.5-32B refer to their respective Instruct models. All training-based models are derived from their corresponding open-source models. The training data of DynaSearcher consist of the stage-2 data from Song et al. (2025) and 8,000 randomly sampled instances from Musique. The RL training epoch is set to 1, with a train batch size of 16 and a learning rate of 1e-6. The KL divergence coefficient set to 1e-3. Each data sample undergoes 8 rollouts during training. We utilize FSDP Zhao et al. (2023) and vLLM333https://github.com/vllm-project/vllm. in VeRL444https://github.com/volcengine/verl. framework, with a sampling temperature of 1.0, top-p of 0.95 and a maximum response length of 8192. For our information gain and penalty reward, we set to 0.5, to 0.9, and to -0.2.
@@ 5 Conclusions
Appendix A Implementation Details
Due to the large size of the dev sets in the 2WikiMultiHopQA and HotpotQA datasets, which affects iteration efficiency, we randomly sample 1,000 examples from their respective dev sets as our final test set, with a fixed random seed 42. We also verify that the performance on this subset is nearly identical to that on the full dev set, indicating that this approach can significantly improve iteration efficiency. To better understand the complexity of multi-hop reasoning in these datasets, we analyze the hop distribution of the HotpotQA, 2WikiMultiHopQA, MuSiQue, MoreHopQA, and Frames dev/test sets in Figure 3. The statistics show that there is a high proportion of complex reasoning queries with 3 hops or more. HotpotQA lacks explicit hop annotations, so we instead count the number of supporting facts.
In the retrieval process, we employ the multilingual-e5-base as the retriever and use the widely used Wikipedia dump from December 2018 as the retrieval corpus, which comprises over 21 million passages. To improve retrieval efficiency, we combine the supporting document passages from five multi-hop datasets with one million randomly sampled documents from the 2018 Wikipedia dump to create our final retrieval corpus.
During the training phase, our training data consist of a total of 8,148 examples from HotpotQA and 2WikiMultiHopQA, which are selected through data selection in R1-Searcher. In addition, we randomly sample 8,000 examples from the training set of MuSiQue to form our final training set. The training consists of 2 epochs, with a train_batch_size of 16 and a learning rate of 1e-6. ppo_mini_batch_size is set to 16. The maximum lengths for prompt and response are set to 512 and 8192. Rollouts are conducted with a batch size of 8 and a temperature of 1.0 to encourage exploration. The KL-divergence regularization coefficient and the clipping ratio are set to 0.001 and 0.2, respectively. All experiments are carried out on eight NVIDIA-H20-96G. In the inference stage, we use SGLang or vLLM as the underlying inference engines and set different maximum context lengths and maximum retrieval times to avoid the impact of outlier samples on training. For the evaluation of other prompt-based baselines, we use the implementations provided in the ReSearch GitHub repository555https://github.com/Agent-RL/ReCall.. For other training-based methods, we evaluate them using their publicly available trained models.
Appendix B Prompt Examples
Although the retrieval process of the knowledge graph (KG) can provide more diverse information for question exploration, it can also introduce noise for the analysis of the initial question. Therefore, we introduce a KG filtering module to control the number of retrieved triples. The specific filtering logic is shown in Table 6. Table 7 presents the implementation prompt for our LLM-as-Judge (LasJ) score. By leveraging larger and more powerful LLMs as judge models, we can achieve more accurate judgments of responses in the multi-hop QA scenarios. This more precise evaluation approach can be incorporated into the training process, which also introduces additional computational overhead for training.
| You are an information extraction expert. Given an input of query, entities and relations, your task is to extract the most relevant evidence(s). If there is no relevant one, return "None". The names of entities or relations can be multiple alias, retain the exact one. For example, |
| <input> {“query”: “Who is the director of Avatar”, “entity”: [“Avatar”], “relation”: [“director”]} </input> <result> |
| {“entity1_id": “x1", “entity1": “Avatar", “relation_id": “y1", “relation": “distinct element of; has class", “entity2_id": “x2", “entity2": “character code; codeset"} {“entity1_id": “x1", “entity1": “Avatar", “relation_id": “y2", “relation": “director; directed", “entity2_id": “x3", “entity2": “James Cameron; james cameron; James’ Cameron"} </result> start extracting: |
| <evidence> entity1; relation; entity2 Avatar; director; James Cameron </evidence> Now, start! |
Appendix C Quantitative Analysis
Table 1 in the main content presents the performance of state-of-the-art LLMs on multi-hop question answering tasks. Interestingly, we find that these models struggle to effectively follow instructions under the search-o1 paradigm, resulting in suboptimal performance. Additionally, in the basic RAG setting, where models are simply asked to answer questions based on retrieved documents, the models tend to respond that no relevant information is found when the answer is not present in the retrieved documents, failing to fully utilize their inherent capabilities. Therefore, after optimizing the prompt for the RAG scenario (see Table 8), the models are able to better integrate their own fundamental abilities with the retrieved information to jointly solve the original questions. As a result, Table 1 in the main text content demonstrates the superior performance of frontier LLMs.
| You will be provided with three pieces of content: the questioner’s question, the user’s response, and the reference answer list. Your task is to score the accuracy of the user’s response based on the criteria outlined below. Please ensure that you carefully read and understand these instructions. |
| Evaluation Criteria: |
| Accuracy - Whether the user’s answer is consistent with the reference answer and answers the questioner’s question. We define this dimension as "whether the user’s response includes all the key points from the reference answer and answers the questioner’s question." |
| Evaluation Steps: |
| 1. Carefully read the questioner’s question and understand its key points. |
| 2. Carefully read the reference answer and understand the key points relevant to the question. |
| 3. Check whether the user’s response includes all the key points from the reference answer and answers the questioner’s question. |
| 4. Based on the evaluation criteria, assign a score in the range of 0 to 5, where 0 indicates that the user’s response does not include any of the key points from the reference answer and completely fails to answer the questioner’s question; 5 indicates that the user’s response includes all the key points from the reference answer and fully and correctly answers the questioner’s question. |
| Example: |
| Questioner’s question: |
| {question} |
| Reference answer: |
| {answer} |
| User’s response: |
| {response} Evaluation result (output only the score between 0 and 5): |
| You are an expert in question answering. Given a question within <question> </question> and some contexts within <context> </context>, you first think about the reasoning process within <think> </think> and put the answer within <answer> </answer>. |
[…]